OpenStudy (anonymous):

subtract 46base 8 from 632base 8 in the octal system

OpenStudy (cwrw238):

this might be a long way but you could convert base 8 to bas 10 and do the subtraction and then convert back to base 8

OpenStudy (cwrw238):

46 base 8 = 6 + 4*8^1 = 38 base 10

OpenStudy (cwrw238):

632 base 8 = 2 + 3*8 + 6* 8^2 = 2 + 24 + 384 = 410 base 10

OpenStudy (cwrw238):

410 - 38 = 372 now you need to convert this to base 8

OpenStudy (anonymous):

Definitely the long way ;)

OpenStudy (cwrw238):

yea - i'm slow today as well!!

OpenStudy (anonymous):

You can set it up just like a base 10 subtraction, just remember that when you borrow, you are borrowing 8, not 10.

OpenStudy (cwrw238):